TWO CONSECUTIVE POSITIVE EVEN INTEGERS ARE SUCH THE SUM OF THEIR SQUARES IS 164. FIND THE TWO NUMBERS.
Answer by mananth(16946)

TWO CONSECUTIVE POSITIVE EVEN INTEGERS ARE SUCH THE SUM OF THEIR SQUARES IS 164. FIND THE TWO NUMBERS.
let one even number be x
the other even number will be (x+2)
x^2+(x+2)^2=164
x^2+x^2+4x+4=164
2x^2+4x-160=0
2(x^2+2x-80)=0
x^2+2x-80=0
x^2+10x-8x-80=0
x(x+10)-8(x+10)= 0
(x+10)(x-8)=0
x=-10 OR x=8
8 & 10 are the positive even integers